Can anyone explain with a simple example, what is meant by substituion method in intergration?

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
We apply substitution method when we have function & its derivative in the integrand.
Example:
I = \int cosx.sin^{2}x.dx
sinx = t
cosx.dx = dt
I = \int t^{2}.dt = \frac{t^{3}}{3} + constant
Thanks & Regards
Jitender Singh
IIT Delhi
askIITians Faculty